# Mobile-Bench-v2: A More Realistic and Comprehensive Benchmark for VLM-based Mobile Agents

### v2

@@ 4.1 Experimental Setup
For Mobile-Bench-v2 common, noisy, and ambiguous splits, we experiment agent frameworks such as AppAgent-v1 and MobileAgent-v2 with different fundamental VLMs: InternVL2-40B (chen2024expanding, ), LLAVA-72B-NEXT (li2024llava, ), Qwen2-VL-72B (wang2024qwen2, ), Llama3.2-90B (dubey2024llama, ), Qwen-VL-Max, GPT-4o (achiam2023gpt, ), and GPT-4v. For mobile agents, we use Cogagent (hong2024cogagent, ), UGround-7B (gou2024navigating, ), OS-Atlas-7B (wu2024atlas, ), UI-Tars (qin2025ui, ), Kimi-VL (team2025kimi, ), and DeepSeek-VL2 (wu2024deepseek, ). To reduce cost, only zero-shot evaluations are done on a subset split of Random-800, which has a similar sub-split distribution to the full split. As shown in Figure 4, the distribution of step and app categories in Random-800 is fully consistent with that of the full dataset, and 800 instructions correspond one-to-one to 800 predefined instruction templates. Simple and complex splits set the maximum number of steps to 20 and 25. For multi-path, the current page and expectation page are provided for self-reflection, while other actions are restricted to the current one. Unlike Appagent, we annotate the widget types using specific numbers (Figure 9).
@@ 4.2 Metrics
We establish four metrics proposed in MobileAgentBench (wang2024mobileagentbench, ) and AUTO-UI (zhang2023you, ),
@@
TYPE: $S_{ttp}/S_{gt}$, where $S_{ttp}$ is the number of predicted actions that match the type of golden actions. For the whole three actions, we use TYPE to check whether the action types are correct.

### v3

@@ 4.1 Experimental Setup
For Mobile-Bench-v2 common, noisy, and ambiguous splits, we experiment agent frameworks such as AppAgent-v1 and MobileAgent-v2 with different fundamental VLMs: InternVL2-40B (Chen et al., 2024), LLAVA-72B-NEXT (Li et al., 2024a), Qwen2-VL-72B (Wang et al., 2024d), Llama3.2-90B (Dubey et al., 2024), Qwen-VL-Max, GPT-4o (Achiam et al., 2023), and GPT-4v. For mobile agents, we use Cogagent (Hong et al., 2024), UGround-7B (Gou et al., 2024), OS-Atlas-7B (Wu et al., 2024b), UI-Tars (Qin et al., 2025), Kimi-VL (Team et al., 2025), and DeepSeek-VL2 (Wu et al., 2024c). To reduce cost, only zero-shot evaluations are done on a subset split of Random-800, which has a similar sub-split distribution to the full split. As shown in Figure 4, the distribution of step and app categories in Random-800 is fully consistent with that of the full dataset, and 800 instructions correspond one-to-one to 800 predefined instruction templates. Simple and complex splits set the maximum number of steps to 20 and 25. For multi-path, the current page and expectation page are provided for self-reflection, while other actions are restricted to the current one. Unlike Appagent, we annotate the widget types using specific numbers (Figure 9).
@@ 4.2 Metrics
We establish four metrics proposed in MobileAgentBench (Wang et al., 2024c) and AUTO-UI (Zhang and Zhang, 2023),
@@
TYPE Accuracy: $S_{ttp}/S_{gt}$, where $S_{ttp}$ is the number of predicted actions that match the type of golden actions. For the whole three actions, we use TYPE to check whether the action types are correct.